You want to mat and frame a 5x7 photograph. The perimeter of the outside of the mat is 44 inches. The mat is twice as wide at the top and bottom as it is at the sides. Find the dimensions of the mat.
Answer by rfer(16322)

width=7.33
length=14.66
-------------
2(2x)+2(x)=44
4x+2x=44
6x=44
x=7.33
--------------
2*7.33+2*14.66=44